### v2

# Wi-Fi Radar via Over-the-Air Referencing: Bridging Wi-Fi Sensing and Bistatic Radar

Abstract
Wi-Fi channel state information (CSI), which is originally acquired for communication purposes, has recently been reused for sensing and radar-like functionalities. However, in practical Wi-Fi systems with independent clocks at the transmitter and receiver, the lack of a common delay and phase reference fundamentally precludes phase-coherent radar-like delay–Doppler analysis. By exploiting the line-of-sight (LoS) path component, i.e., the earliest-arriving direct path, as an over-the-air (OTA) reference for delay and phase, we propose an OTA LoS-path referencing scheme, termed LoSRef, that enables delay calibration and phase alignment under this practical constraint. Unlike conventional Wi-Fi bistatic radar systems that rely on wired reference signals or dedicated reference antennas, the proposed LoSRef-based framework enables phase-coherent bistatic radar-like operation that can be integrated into typically deployed Wi-Fi systems. Through human gait and respiration experiments in indoor environments, we demonstrate that phase-coherent channel impulse responses and corresponding delay–Doppler responses can be obtained using only commodity Wi-Fi devices. This enables physically interpretable human motion sensing, including gait-induced range variation and respiration-induced sub-wavelength displacement, as well as the extraction of target-induced dynamics up to 20 dB weaker than dominant static multipath components.
I Introduction
As a key application of integrated sensing and communications (ISAC), which aims to augment wireless communication systems with sensing capabilities by leveraging existing wireless infrastructure [1], wireless LAN systems with device-free sensing capability have attracted significant attention in recent years [2, 3, 4]. By exploiting ubiquitous Wi-Fi transmissions, numerous studies have demonstrated the feasibility of monitoring human respiration, gait, and daily activities without requiring the target to carry any dedicated device. These capabilities have enabled a broad range of applications in healthcare, smart living environments, and ambient intelligence, while benefiting from the low deployment cost and widespread availability of commercial Wi-Fi networks.
Early Wi-Fi sensing studies primarily relied on coarse metrics such as received signal strength indicator (RSSI) [5, 6]. A major turning point was the release of the Linux 802.11n channel state information (CSI) tool [7], which enabled researchers to access CSI from commercial off-the-shelf (COTS) IEEE 802.11n devices. In this context, CSI refers to the complex-valued channel frequency response (CFR) measured across orthogonal frequency division multiplexing (OFDM) subcarriers. From a communication-system perspective, CSI is obtained for communication purposes such as equalization, precoding, and link adaptation, rather than being designed specifically for sensing purposes. Throughout this paper, the terms CSI and CFR are used interchangeably: CSI is used when referring to channel measurement and acquisition, while CFR denotes the frequency-domain representation of the channel and its Fourier-transform relationship to the channel impulse response (CIR), which represents the channel in the delay domain.
This CSI acquisition capability has triggered extensive research on CSI-based Wi-Fi sensing, including motion-related sensing based on temporal variations of the wireless channel, as summarized in representative surveys such as [2, 3, 4]. More recently, Wi-Fi sensing has matured beyond ad-hoc research prototypes, culminating in the standardization of IEEE 802.11bf [8], which defines channel measurement procedures to explicitly support a wide range of sensing use cases, including range, motion, and gesture detection, within Wi-Fi networks.
Figure 1: Comparison of reference acquisition strategies. While (a) and (b) rely on dedicated reference hardware, the proposed OTA LoS-path referencing scheme (LoSRef) in (c) exploits the Tx–Rx LoS path as a delay reference, eliminating the need for auxiliary reference hardware. Moreover, (c) is compatible with a single-port Rx, whereas (a) and (b) require multi-port Rxs.
Alongside the aforementioned Wi-Fi sensing approaches, there also exists a line of research that exploits COTS Wi-Fi transmissions for radar applications [9, 10]. For clarity, we refer to methods operating without explicit reference signals as Wi-Fi sensing, and those employing explicit reference signals for phase-coherent analysis as Wi-Fi radar. Early Wi-Fi radar studies typically relied on measurement-grade receivers and external reference acquisition mechanisms, such as reference antennas or wired signal taps, as illustrated in Fig. 1(a) and (b), to enable coherent delay or Doppler estimation [9, 11, 10, 12, 13]. More recent works have demonstrated that reference signals can also be exploited in fully COTS-based systems to stabilize CSI measurements, i.e., the phase of the underlying CFR [14, 15, 16, 17]. These reference-aided architectures are fundamentally consistent with bistatic radar principles and clearly demonstrate the potential of Wi-Fi signals for radar-like sensing. However, due to their reliance on cable routing or dedicated reference antenna placement, most prior Wi-Fi radar functionalities cannot be integrated into typically deployed Wi-Fi systems.
These observations reveal a fundamental gap between existing Wi-Fi sensing and Wi-Fi radar approaches. While reference-aided Wi-Fi radar enables physically meaningful delay and Doppler analysis, such architectures cannot be integrated into typical Wi-Fi systems. Conversely, Wi-Fi sensing approaches that rely, for example, on CFR magnitude alone do not require any modification to existing Wi-Fi systems, but lack direct access to physically interpretable propagation parameters.
This raises a key question: Can bistatic Wi-Fi radar functionality be realized using only unmodified COTS Wi-Fi devices, without wired references or dedicated reference antennas, while being integrated into typical communication-oriented Wi-Fi systems, and still enabling physically meaningful delay–Doppler estimation?
The contributions of this paper are summarized as follows.
  • Unlike most existing Wi-Fi sensing approaches that operate directly on the CFR, which is originally acquired for communication purposes, a CIR is constructed using CFR obtained from COTS IEEE 802.11ax devices with $160\text{\,}\mathrm{MHz}$ bandwidth [18]. This is enabled by carefully designed frequency-domain preprocessing of the raw CFR, and provides a reliable basis for subsequent phase-coherent delay-domain processing.
  • To the best of our knowledge, this is the first work that realizes reference-aided bistatic Wi-Fi radar that can be integrated into typically deployed Wi-Fi systems, using only unmodified COTS Wi-Fi devices, without relying on wired references or dedicated reference antennas. To this end, we propose an over-the-air (OTA) line-of-sight (LoS)-path referencing scheme (LoSRef) that exploits the temporal stability of the transmitter–receiver (Tx–Rx) LoS path to provide a common delay and phase reference in practical Wi-Fi systems, as illustrated in Fig. 1(c), thereby enabling phase-coherent CIR acquisition.
  • The proposed LoSRef-based framework requires only a single antenna port at the Rx, whereas conventional reference-aided bistatic Wi-Fi radar functionalities typically rely on multi-port Rxs or dedicated reference channels. This further simplifies hardware requirements and facilitates integration into typical Wi-Fi systems using unmodified COTS Wi-Fi devices.
  • By leveraging the phase of the CIR, we demonstrate that respiration-induced body-surface motion, as a representative example of sub-wavelength displacement, can be directly observed and quantitatively interpreted in indoor environments, in a manner that is not accessible with conventional magnitude-based Wi-Fi sensing.
  • We experimentally show that a delay–Doppler representation can be derived from the CIR while preserving the sign of the Doppler frequency, which is inherently lost in magnitude-based approaches. This enables discrimination between approaching and receding motions, achieving phase-coherent bistatic radar functionality with COTS Wi-Fi devices.
While the proposed framework enables various application-level sensing tasks, including gait analysis and respiration monitoring, the focus of this paper is on establishing a delay-domain, phase-coherent signal representation enabled by LoSRef. Accordingly, the intent is not to compete with state-of-the-art application-specific estimators or to benchmark their estimation accuracy, but to establish a physically grounded, phase-coherent signal representation that fundamentally differs from conventional magnitude-based Wi-Fi sensing and upon which such estimators and applications can be built.
This paper is organized as follows. Section II reviews related work on Doppler-based Wi-Fi sensing and reference-aided Wi-Fi radar. Section III discusses the reference ambiguity due to independent clocks in practical Wi-Fi systems. Section IV presents the proposed phase-coherent CIR construction framework, including delay and phase referencing based on a stable LoS-path reference. Section V demonstrates bistatic range estimation through human gait and respiration experiments. Section VI presents delay–Doppler response estimation results for moving targets. Section VII provides additional discussion, followed by Section VIII, which concludes the paper.
II Related Work
We review prior Wi-Fi sensing studies relevant to this work, followed by reference-aided Wi-Fi radar approaches. Estimation of propagation distance or time-of-flight based on statistical or geometry-based features of CSI [19, 20] is outside the scope of this review, as such studies primarily aim at inferring static distance or position from channel statistics, rather than resolving delay–Doppler responses from phase-coherent CIRs.
II-A Doppler-Based Wi-Fi Sensing
This subsection focuses on Wi-Fi sensing studies that primarily perform Doppler-related spectral analysis without explicit reference signals, which are most closely related to the present work. In these approaches, spectral analysis techniques such as the short-time Fourier transform (STFT) are typically applied to the magnitude of the CFR or its low-dimensional representations. This design choice is mainly due to the fact that the phase of the CFR is generally unstable across packet acquisitions on COTS Wi-Fi devices, making it difficult to reliably exploit phase information over time. Although antenna-wise or subcarrier-wise phase differences have been successfully exploited for localization [21, 22], such phase-based quantities are not typically treated as time-domain signals for subsequent time–frequency spectral analysis.
In [23], respiration sensing is performed by selecting subcarriers whose CFR magnitudes exhibit strong target-induced periodic variations. Spectral analysis is then applied to the selected magnitude-domain time series to estimate the respiration rate.
WiFiU [24] adopts a principal component analysis (PCA)-based dimensionality reduction of the CFR magnitude, followed by STFT on the resulting PC scores to extract Doppler-related information. Widar [25] further interprets the dominant spectral components as path length change rates, and estimates target velocity and location through an explicit geometric model. It is worth noting that, in both approaches, the extracted Doppler-related quantities are derived from magnitude-domain spectral analysis.
In particular, Li et al. [26] provide an insightful analysis closely related to the lack of stable phase reference addressed in this work. The authors contrast Wi-Fi sensing and Wi-Fi radar by clarifying what physical quantities are actually estimated in each case. They show that magnitude-based CFR processing inherently loses Doppler sign information and physically meaningful phase evolution because magnitude-only processing removes the phase evolution induced by propagation path-length changes. As a result, the extracted spectral components reflect periodic signal strength fluctuations rather than a true Doppler frequency defined by wave propagation.
TABLE I: Positioning of This Work with Respect to Existing Studies Including ISAC-Enabling Capability
Study Tx Rx Reference Signal (Fig. 1) ISAC-enabling Standard
Typical Wi-Fi sensing [24, 23, 25] COTS COTS $\times$ IEEE 802.11n
[10, 12, 13] COTS Instrument Rx (a) Wired reference $\times$ IEEE 802.11g
[9, 11] COTS Instrument Rx (b) Reference antenna $\times$ IEEE 802.11g
[14, 16] COTS COTS (a) Wired reference $\times$ IEEE 802.11n
[15] COTS COTS (b) Reference antenna $\times$ IEEE 802.11n
[17] COTS COTS (b) Reference antenna $\times$ IEEE 802.11ax
This work COTS COTS (c) LoSRef: OTA LoS-path referencing IEEE 802.11ax
II-B Reference-Aided Wi-Fi Radar
Table I summarizes the positioning of this work with respect to representative Wi-Fi sensing and reference-aided Wi-Fi radar studies, highlighting differences in Tx and Rx types, reference signals, ISAC-enabling capability, and supported standards. Here, ISAC-enabling indicates whether such functionality can be realized within typically deployed Wi-Fi systems without additional reference hardware. In [9, 11], a bistatic radar architecture is adopted, where a COTS Wi-Fi Tx and a measurement-grade Rx are used. The Rx employs separate reference and surveillance antennas, as illustrated in Fig. 1(b), with the antenna configuration designed such that target-induced variations are primarily captured in the Tx–surveillance path. The target-induced Doppler frequency is then extracted by correlating the reference and surveillance signals.
As an alternative reference acquisition strategy, [10, 12, 13] inserted a directional coupler between a COTS Wi-Fi Tx and its antenna to obtain a wired reference signal, as illustrated in Fig. 1(a). It should be noted that in all these studies [9, 11, 10, 12, 13], the Rx is instrument-grade, and sensing is performed directly on received signals rather than on CSI measurements obtained from COTS Wi-Fi devices. Moreover, these studies largely predate the widespread availability of CSI on COTS Wi-Fi hardware, which emerged later with IEEE 802.11n and CSI extraction tools [7]. In this sense, Wi-Fi is exploited merely as a signal of opportunity, without leveraging the CSI natively observed by COTS Wi-Fi devices. Nevertheless, these works clearly demonstrate that Wi-Fi signals can fundamentally support radar functionality.
Building on these reference-based architectures, several studies have demonstrated fully COTS-based implementations, in which both the Tx and Rx are COTS Wi-Fi devices, while still exploiting additional reference hardware for phase calibration. In [14, 16], a wired reference is employed. This enables explicit Doppler frequency estimation for fine-grained hand gesture recognition [14] and respiration monitoring in dynamic environments [16].
Similarly, WiEps [15] and SoundiFi [17] adopt a reference antenna to achieve phase-stable CSI measurements under a fully COTS-based setup. These capabilities are exploited for dielectric property measurement of materials [15] and for channel sounding [17], respectively. These results indicate that wired or antenna-based reference signals remain an effective and reliable means of phase-coherent radar-like analysis using Wi-Fi devices. Note that although not all of these studies explicitly use the term “radar,” their architectures are consistent with bistatic radar principles, as they rely on explicit reference signals to enable phase-coherent analysis.
In contrast to existing reference-aided Wi-Fi radar systems, which rely on wired references or dedicated reference antennas, the LoSRef exploits the OTA Tx–Rx LoS path. As a result, phase-coherent bistatic radar operation is achieved using only a single-antenna, single-port Rx, without requiring multi-port Rx architectures, thereby enabling integration into typical Wi-Fi systems and facilitating ISAC operation.
III CIR Representation and Reference Ambiguity Due to Free-Running Clocks
This section briefly reviews the CIR representation and highlights the reference ambiguity arising from free-running clocks in practical Wi-Fi systems, which fundamentally prevents phase-coherent analysis and physically meaningful time-domain processing.
III-A CIR Representation and Doubly Selective Channels
The equivalent low-pass representation of the CIR can be described using a tapped delay line (TDL) model [27]. Under this model, the CIR at a given measurement time $t$ is expressed as a superposition of discrete propagation paths, each characterized by a complex path gain and a propagation delay, as
$h(\tau,t)=\sum_{\ell=1}^{L}\eta_{\ell}\,\mathrm{e}^{-\mathrm{j}2\pi f_{\mathrm{c}}\tau_{\ell}}\delta(\tau-\tau_{\ell}),$ (1)
where $L$ denotes the number of discrete propagation paths, $\eta_{\ell}$ and $\tau_{\ell}$ represent the complex amplitude and delay of the $\ell$-th path, respectively, and $f_{\mathrm{c}}$ is the carrier frequency. Here, $\tau$ denotes the signal propagation delay measured from the transmission time, and $t$ denotes the observation time of the CIR, corresponding to successive channel measurements rather than a continuous-time variable. This representation captures the multipath structure of the wireless channel in the delay domain, where the complex phase of each path directly reflects the propagation path length.
In dynamic environments, the underlying wireless channel is doubly selective, exhibiting variations across both delay and time. The CIR $h(\tau,t)$ provides a delay–time representation that captures such joint delay–time variations.
III-B Reference Ambiguity Due to Free-Running Clocks
In Wi-Fi systems, where the Tx and Rx employ independent, free-running clocks, neither a common timing reference nor an absolute phase reference is available between them. As a result, the CIR observed at the Rx does not preserve the absolute delay reference to the transmission timing, nor does it retain a phase-coherent representation across successive observations. Accordingly, the observed CIR $h_{\mathrm{raw}}(\tau,t)$ is related to the underlying physical CIR $h(\tau,t)$ through an unknown delay offset and an unknown phase offset, and is given by
$h_{\mathrm{raw}}(\tau,t)=h(\tau+\alpha_{t},t)\,\mathrm{e}^{\mathrm{j}\beta_{t}},$ (2)
where $\alpha_{t}$ and $\beta_{t}$ denote the delay offset and phase offset, respectively, which may vary across observation times $t$ [21, 22]. These offsets originate from free-running clocks at the Tx and Rx, and prevent direct use of absolute delay and phase information, thereby precluding phase-coherent analysis of the doubly selective channel in practical Wi-Fi systems without a shared reference. Note that the delay offset $\alpha_{t}$ of the devices used in this study has been experimentally characterized in [17].
When (2) is viewed in the frequency domain, the resulting complex-valued CFR $H_{\mathrm{raw}}(f,t)$, where $f$ denotes the OFDM subcarrier frequency, is inherently discontinuous with respect to time $t$ due to these unknown offsets, which precludes direct temporal or spectral analysis. To circumvent this issue, many Wi-Fi sensing studies, including [23], operate on the magnitude of the CFR, $|H_{\mathrm{raw}}(f,t)|$. Taking the magnitude effectively removes these unknown offsets, yielding an apparently smooth time series that is amenable to spectral analysis. This convenience, however, comes at the cost of discarding the phase evolution associated with wave propagation, thereby precluding any phase-coherent interpretation of propagation delay or Doppler frequency.
In contrast, in bistatic radar systems where a reference signal is available, a phase-coherent CIR $h(\tau,t)$ can be recovered. This temporal continuity enables the use of the Fourier-transform duality between $h(\tau,t)$ and physically interpretable delay–Doppler representations [28, 27].
Measured channel frequency response (CFR)Frequency-domain preprocessingOutlier removal; DC phase interpolation;Magnitude equalizationRaw CFR $H_{\mathrm{raw}}(f,t)$IDFT (Blackman window, zero-padding)Raw channel impulse response (CIR) $h_{\mathrm{raw}}(\tau,t)$LoSRef: OTA LoS-path referencingPeak detectionDelay calibration & phase alignmentPhase-coherent CIR $h(\tau,t)$Clutter removalResidual CIR $\tilde{h}(\tau,t)=h(\tau,t)-\mathbb{E}_{t}[h(\tau,t)]$Time-domain processing for DopplerUniform resampling ($\varDelta t$)STFT (Hann window, zero-padding, mean removal)Delay–Doppler response $s(\tau,\nu;t)$
Figure 2: Processing flow from measured CFR to the delay–Doppler response $s(\tau,\nu;t)$. The LoSRef exploits the Tx–Rx line-of-sight path as a physical delay and phase reference, enabling phase-coherent alignment of CIR across packet acquisitions.
IV Phase-Coherent Time-Variant Impulse Response Estimation
The offsets $\alpha_{t}$ and $\beta_{t}$ in (2) are unknown and vary across observation times, which prevents phase-coherent alignment of CIRs across successive packet acquisitions. In contrast, the delay of the LoS-path component in (1) is determined solely by the Tx–Rx distance and is therefore expected to remain constant as long as the Tx and Rx are stationary. Exploiting this temporal stability, the LoSRef uses the LoS-path component as an OTA reference to mitigate the effects of these unknown offsets and to enable phase-coherent time-variant CIR estimation. Fig. 2 summarizes the overall processing flow of the proposed framework, incorporating LoSRef, from the measured CFR to the delay–Doppler response, and each processing block is detailed in the following subsections.
IV-A CIR Construction for Robust LoS-Path Identification
Recall that $H_{\mathrm{raw}}(f,t)$ denotes the CFR observed at time $t$. The corresponding CIR is obtained by transforming the CFR from the frequency domain to the delay domain via an inverse Fourier transform, given by
$h_{\mathrm{raw}}(\tau,t)=\mathcal{F}^{-1}[H_{\mathrm{raw}}(f,t)].$ (3)
In practice, the CFR is further processed by applying windowing and effective bandwidth extension, as described below. Strictly speaking, the term $H_{\mathrm{raw}}(f,t)$ in (3) should therefore be interpreted as the preprocessed CFR.
It should be noted that the resulting CIR includes the phase and gain of the transceivers and the radio channel, and is convolved with the impulse response of the applied frequency-domain window. For simplicity, this windowed version is hereafter referred to as the CIR. Similarly, although the observations are discrete and the transformation should formally be written as an inverse discrete Fourier transform (IDFT), we retain the simplified representation in (3) to avoid unnecessary notational complexity.
To construct a stable delay-domain CIR, frequency-domain windowing and delay-domain oversampling are applied, following the same principle as in our prior Wi-Fi channel sounding work [17, Section III-C], with window and oversampling parameters adapted to the present setup. Specifically, a Blackman window is used prior to the inverse Fourier transform to suppress delay-domain sidelobes. In addition, delay-domain oversampling with a factor of $\kappa=32$ is performed via zero padding in the frequency domain. It is emphasized that such oversampling does not improve the intrinsic delay resolution, which remains fundamentally limited by the signal bandwidth ($1/B=$6.3\text{\,}\mathrm{ns}$\$), but is nevertheless essential for achieving accurate and temporally stable peak detection under discretized delay sampling.
IV-B LoSRef-Based Delay Calibration and Phase Alignment
Let the known LoS OTA Tx–Rx propagation distance $d_{\text{Tx--Rx}}$ be used as the reference distance $d_{\mathrm{ref}}$. For clarity, this OTA LoS-path referencing scheme is hereafter simply referred to as LoSRef. While the measurement accuracy of $d_{\mathrm{ref}}$ may warrant careful discussion depending on the target application, this issue is beyond the scope of this work, since $d_{\mathrm{ref}}$ is used solely for absolute delay-axis alignment. Doppler frequency estimation, which is based on the temporal derivative of the propagation phase, is therefore independent of the absolute value of $d_{\mathrm{ref}}$, as long as it is constant over time.
The propagation delay of the LoS path relative to the transmitted signal is given by $\tau_{\mathrm{ref}}=d_{\mathrm{ref}}/c$, where $c$ denotes the speed of light. In an ideal delay-synchronized system, the LoS-path component would therefore appear at $\tau_{\mathrm{ref}}$ in the CIR.
In practice, however, this alignment is not observed, because the Tx and Rx are not delay-synchronized and the delay axis of the measured CIR is not referenced to the actual packet transmission timing. As a result, the absolute delay origin varies across packet acquisitions.
To address this issue, the delay axis of the CIR is calibrated by identifying the dominant component, assumed to correspond to the LoS path, and aligning its delay with the reference delay $\tau_{\mathrm{ref}}$. This procedure establishes a common delay reference across packet acquisitions. [^1] Strictly speaking, the identified peak corresponds to a delay bin dominated by the LoS path rather than an isolated propagation path. Specifically, the delay calibration is performed to satisfy
$\operatorname*{arg\,max}_{\tau}\,\lvert h(\tau,t)\rvert=\tau_{\mathrm{ref}}.$ (4)
Figure 3: LoSRef: OTA LoS-path referencing for delay calibration and phase alignment. The dominant LoS-path component is used as an OTA reference to compensate unknown offsets arising from free-running clocks at the Tx and Rx.
This calibration is carried out as follows. For each time index $t$, the delay and phase of the strongest signal in the raw CIR $h_{\mathrm{raw}}(\tau,t)$ are first identified as
$\tau_{t}^{\star}$ $=\operatorname*{arg\,max}_{\tau}\,\lvert h_{\mathrm{raw}}(\tau,t)\rvert$, (5)
$\theta_{t}^{\star}$ $=\arg\,h_{\mathrm{raw}}(\tau_{t}^{\star},t).$ (6)
The calibrated CIR is then obtained by compensating for both the delay offset and the phase rotation of this strongest signal, which jointly restores phase coherence across successive CIR measurements, given by
$h(\tau,t)=h_{\mathrm{raw}}(\tau-(\tau_{\mathrm{ref}}-\tau_{t}^{\star}),t)\,{\rm e}^{-{\rm j}\theta_{t}^{\star}},$ (7)
as illustrated in Fig. 3. The above calibration procedure follows the same mathematical formulation as our prior channel sounding work [17].
In addition, the magnitude of the CIR is normalized such that $\lvert h(\tau_{\mathrm{ref}},t)\rvert$ matches the free-space path gain corresponding to the reference distance $d_{\mathrm{ref}}$. Consequently, the resulting CIR $h(\tau,t)$ represents the complex path gain associated with each propagation path.
In this calibration procedure, the delay shift by $\tau_{\mathrm{ref}}-\tau_{t}^{\star}$ is referred to as delay calibration, while the phase rotation by ${\rm e}^{-{\rm j}\theta_{t}^{\star}}$ is referred to as phase alignment. Through phase alignment, the phase of each multipath component is expressed relative to that of the dominant (LoS-path) component. As a result, the phase of the CIR becomes continuous across successive acquisitions, enabling phase-coherent time-variant and Doppler analysis.
The LoSRef-based calibration relies on the assumption that the LoS-path component remains dominant and is well resolvable in the CIR. Scenarios in which this assumption is violated, such as extremely close target proximity to the Tx–Rx pair, are beyond the scope of this work.
IV-C Clutter Removal via LoSRef-based CIR Alignment
As in (1), the observed CIR $h(\tau,t)$ can be modeled as a superposition of static multipath components, hereafter referred to as clutter, and time-varying components induced by target motion. In indoor environments, the CIR is typically dominated by strong clutter, which tends to mask relatively weak motion-induced variations associated with human targets.
To isolate the time-varying component, clutter is removed using a simple temporal averaging approach. Specifically, the clutter is estimated by taking the temporal average of the CIR over the observation interval, i.e.,
$\overline{h}(\tau)\coloneqq\mathbb{E}_{t}[h(\tau,t)].$ (8)
The time-varying component is then extracted by subtracting the estimated clutter from the original CIR:
$\tilde{h}(\tau,t)\coloneqq h(\tau,t)-\overline{h}(\tau).$ (9)
The resulting signal, referred to as the residual CIR $\tilde{h}(\tau,t)$, is used for subsequent sensing and analysis.
In radar signal processing, clutter removal is a standard operation in which static components, including the direct path, are suppressed. In contrast, the proposed LoSRef-based framework differs in that the LoS-path component is first exploited as a physical delay and phase reference to achieve CIR alignment, and subsequently suppressed through clutter removal.
Note that, in conventional Wi-Fi sensing systems, direct processing of complex-valued CIRs is not physically meaningful due to the lack of a shared delay and phase reference between the Tx and Rx, which prevents meaningful temporal averaging. For example, Widar [25] extracts motion-related components by applying PCA to the CFR magnitude, followed by time–frequency spectral analysis. By contrast, the proposed LoSRef-based approach aligns CIRs on a common delay axis using a stable LoS-path reference, enabling phase-coherent temporal averaging and delay-domain clutter removal, which may be advantageous for slowly varying motions.
V Experimental Analysis of OTA Phase-Referenced Channel Impulse Responses
To validate the delay-domain sensing framework described in Section IV, we conducted two representative experimental scenarios designed to capture both (i) coarse delay-domain magnitude variations induced by translational human gait and (ii) sub-wavelength, phase-sensitive variations induced by respiration.
V-A Experimental Equipment
Figure 4: A COTS Wi-Fi device used as the Tx or Rx.
As the Tx and Rx, we used two ASUS NUC 13 Rugged-Tall devices, shown in Fig. 4, each equipped with an Intel AX210 Wi-Fi card and running PicoScenes [29] for CSI acquisition. Only a single external antenna was used at each device, resulting in a single-input single-output channel. In the following, the Tx and Rx positions are defined by the locations of the corresponding antennas. The devices were connected to a control PC via Ethernet and remotely operated to coordinate packet transmission and reception.
TABLE II: Transmission and CIR Construction Parameters
Standard IEEE 802.11ax [18]
Number of Tx/Rx antennas (used) 1 / 1
Center frequency $f_{\mathrm{c}}$ $5570\text{\,}\mathrm{MHz}$
Wavelength $\lambda$ $53.8\text{\,}\mathrm{mm}$
Bandwidth $B$ $160\text{\,}\mathrm{MHz}$
Subcarrier spacing $\varDelta f$ $78.125\text{\,}\mathrm{kHz}$
Number of observed subcarriers $1992$
Number of subcarriers after interpolation $2025$
Subcarrier indices 1012-1012, 1011-1011, …, $1012$
Nominal packet transmission interval $\varDelta t$ $1\text{\,}\mathrm{ms}$
Number of FFT points w/o interpolation $2048$
Delay-domain interpolation rate $\kappa$ $32$
Frequency-domain window Blackman
The Tx transmitted packets using the parameters summarized in Table II. The Rx received these packets and acquired the corresponding CSI, along with the associated RSSI. The nominal packet transmission interval was set to $1\text{\,}\mathrm{ms}$; however, the actual inter-packet intervals observed at the Rx are not perfectly uniform. The characteristics of this timing irregularity are discussed in detail in Section VI.
For ground-truth position reference in the human gait experiment, HTC VIVE Tracker 3.0 [30] based on infrared optical tracking, together with SteamVR Base Station 2.0, were attached to the top of the moving human target’s head and to the Tx and Rx.
V-B Frequency-Domain Preprocessing
In addition to the standard preprocessing provided by the PicoScenes Python Toolbox [29], such as interpolation of unobserved pilot subcarriers, further frequency-domain preprocessing steps were applied, partially following the procedure described in Section VI-C of [17], with additional preprocessing steps introduced in this study. After applying these preprocessing steps, the resulting frequency-domain channel representation is denoted as $H_{\mathrm{raw}}(f,t)$. The details of the preprocessing procedure are as follows.
As an outlier removal procedure, CFR samples that exhibit abrupt and extremely low RSSI are treated as outliers and excluded from further analysis. In the considered bistatic configuration, target motion affects the CFR, but is not expected to induce large changes in the overall received power, which is dominated by the LoS path. Accordingly, sudden power drops exceeding $10\text{\,}\mathrm{dB}$ are regarded as non-physical outliers since the received RSSI, reported with $1\text{\,}\mathrm{dB}$ granularity, typically varies only on the order of a few decibels.
Figure 5: Impact of phase and power preprocessing on the CFR.
Phase interpolation [17] was applied to the DC subcarriers (indices 11-11 to $11$), which carry no transmitted signal in IEEE 802.11ax [18]. As shown in Fig. 5LABEL:sub@fig:cfr_phase, prior to this correction, ignoring the inherent $2\pi$ phase ambiguity results in anomalous inter-subcarrier phase differences at the DC subcarriers. The inter-subcarrier phase difference is defined as
[…]
where $k\in\{$-1012$,$-1011$,\ldots,$1012$\$\} denotes the subcarrier index, with the corresponding frequency given by $f=k\varDelta f$ and the subcarrier spacing $\varDelta f=$78.125\text{\,}\mathrm{kHz}$\$.
To resolve this issue, the phase interpolation estimates the phase trend from neighboring subcarriers using the unwrapped phase. The DC subcarrier phases are then reconstructed accordingly, as confirmed by the resulting phase differences in the DC region being consistent with those of the surrounding subcarriers.
As shown in Fig. 5LABEL:sub@fig:cfr_magnitude, pronounced attenuation appears around frequency offsets of approximately $\pm$60\text{\,}\mathrm{MHz}$\$. Following the same treatment as in [17], this attenuation is considered to originate from device-specific characteristics rather than wave propagation effects. Accordingly, the CFR values at the corresponding subcarriers are excluded, and their magnitudes are interpolated from those of the neighboring subcarriers.
Figure 6: Measured and compensated CFR magnitudes for a single snapshot with direct coaxial connection, with the estimated attenuation model overlaid.
In addition, as observed in Fig. 5LABEL:sub@fig:cfr_magnitude, the CFR magnitude exhibits noticeable attenuation for frequencies with $|f|>$53\text{\,}\mathrm{MHz}$\$. To quantitatively characterize this attenuation, the Tx and Rx were directly connected using coaxial cables and an attenuator, thereby forming a non-frequency-selective channel. Fig. 6 shows an example of the measured CFR, where the magnitude is normalized by the RMS value over subcarriers with $|k|\leq$680$\$. The measured CFR magnitude is approximately unity for subcarriers with $|k|\leq$680$\$, while it decreases to about $0.6$ for $|k|\geq$704$\$, with a gradual and nearly linear transition observed between these regions. To quantitatively capture this attenuation behavior, we introduce a simple piecewise attenuation model $A(k)$ in this study as
$A(k)=\begin{cases}1,&|k|\leq$680$,\\ -(|k|-$680$)/60+1,&$680$<|k|<$704$,\\ $0.6$,&$704$\leq|k|,\end{cases}$ (11)
which approximates the observed magnitude trend and is overlaid in Fig. 6.
Based on this model, the non-uniform frequency response is compensated by scaling the magnitude of each subcarrier by the inverse of $A(k)$. This operation effectively equalizes the spectral magnitude, resulting in a flattened and approximately uniform CFR across all subcarriers, as shown in Fig. 6.
We confirmed consistent frequency-domain attenuation and phase behavior across three or more ASUS NUC devices. Notably, different device sets were used in Sections V-D and V-E, yet comparable characteristics were observed. This suggests that device-to-device variability is limited for the hardware used in this study, and that the proposed preprocessing is practically applicable under similar experimental conditions.
V-C Impact of Preprocessing on CIR Construction
Figure 7: Impact of phase and power preprocessing on the CIR. Note that the horizontal axis of the “Without oversampling” curve is manually shifted to match that of the “Before phase/power preprocessing” curve for visualization only.
Fig. 7 examines the impact of preprocessing on the constructed CIR. The delay axis is referenced to the strongest path, which is assumed to correspond to the LoS path. In the LoS Tx–Rx experimental environment considered here, negative delays correspond to non-physical artifacts. The figure also illustrates how delay-domain oversampling helps alleviate discretization effects in peak delay estimation.
Before phase/power preprocessing, a noticeable pre-LoS component was observed at approximately $26$\,\mathrm{dB}$, exceeding the expected Blackman-window sidelobe level of about $58$\,\mathrm{dB}$. After preprocessing, such components were suppressed to the sidelobe level, indicating effective mitigation of spurious responses. Minor residual discrepancies may still remain due to non-ideal hardware effects.
The wide bandwidth of $160\text{\,}\mathrm{MHz}$ available in IEEE 802.11ax yields a well-localized impulse response with suppressed sidelobes, providing a stable and temporally consistent LoS-path peak. From an estimation-theoretic perspective, a wider bandwidth also tightens the Cramér–Rao lower bound for delay estimation [31], thereby improving the achievable delay estimation accuracy. This contrasts with many existing Wi-Fi sensing studies based on $20\text{\,}\mathrm{MHz}$ or $40\text{\,}\mathrm{MHz}$ channels of IEEE 802.11n devices, where limited bandwidth results in a less distinct delay-domain structure. Consequently, the LoS-path component serves as a reliable physical reference for LoSRef and subsequent phase-coherent delay-domain analysis.
V-D Range Estimation of Translational Human Gait
V-D1 Experimental Setup and Range–Time Representation
Figure 8: Experimental setup for the human gait experiment.
The Tx and Rx were placed $1.0\text{\,}\mathrm{m}$ apart in a LoS configuration, corresponding to a reference delay of $\tau_{\mathrm{ref}}=$1.0\text{\,}\mathrm{m}$/c=$3.3\text{\,}\mathrm{ns}$\$, with both antennas positioned at a height of approximately $0.8\text{\,}\mathrm{m}$ above the floor. The experimental setup and the straight-line walking trajectory representing back-and-forth translational gait motion of the human target were illustrated in Fig. 8. This simple gait scenario, involving back-and-forth translational motion along a straight-line trajectory, allows the time-varying CIR and its Doppler characteristics to be examined in a clear and interpretable manner, which will be discussed later in Section VI.
Figure 9: Snapshot of the observed CIR $h(\tau,t)$, the estimated clutter $\overline{h}(\tau)$, and the resulting residual CIR $\tilde{h}(\tau,t)$ at $t=$3.2\text{\,}\mathrm{s}$\$, where $h(\tau,t)$ and $\tilde{h}(\tau,t)$ are averaged over 10 measurements.
Figure 10: Heatmap representations of the CIR in the delay–time domain measured in the presence of a moving human target. Note that different color scale limits are used in (a) and (b), as the residual CIR in (b) has significantly lower power than the observed CIR in (a).
Fig. 9 shows a snapshot of the calibrated CIR $h(\tau,t)$ at a representative time instant $t=$3.2\text{\,}\mathrm{s}$\$, whereas Fig. 10LABEL:sub@fig:h1_tau-t presents the corresponding CIR evolution over the entire observation interval as a delay–time heatmap. For visualization purposes, the delay axis $\tau$ is converted to propagation distance $c\tau$, yielding a range–time map. In the bistatic configuration considered here, this propagation distance represents the sum of the Tx–target and target–Rx path lengths. Following bistatic radar terminology [32], this quantity is hereafter referred to as the bistatic range $d_{\mathrm{Tx\text{--}target\text{--}Rx}}$.
Fig. 10LABEL:sub@fig:h1_tau-t shows that the strongest signal was consistently aligned at $\tau_{\mathrm{ref}}=$3.3\text{\,}\mathrm{ns}$\$ over time as a result of the LoSRef-enabled delay calibration. Moreover, the time-varying components induced by human motion are not clearly distinguishable in the raw CIR representation.
Accordingly, clutter removal described in Section IV-C is applied to obtain the residual CIR. Fig. 10LABEL:sub@fig:h_tau-t shows the resulting delay–time response, with the geometry-based bistatic range $d_{\mathrm{Tx\text{--}target\text{--}Rx}}$ overlaid for reference. A good agreement was observed between the two.
In addition, a snapshot of the residual CIR and the estimated clutter $\overline{h}(\tau)$ are shown in Fig. 9, together with the previously presented results. Although the human-induced scattering component around $\tau=$25\text{\,}\mathrm{ns}$\$ is about $20\text{\,}\mathrm{dB}$ weaker than the dominant clutter component $\overline{h}(\tau)$, the residual CIR $\tilde{h}(\tau,t)$ clearly captured the corresponding temporal variations through clutter removal.
V-D2 Short-Time Phase Evolution and Doppler Interpretation
Figure 11: Phase of residual CIR $\tilde{h}($20\text{\,}\mathrm{ns}$,t)$ showing continuous phase evolution induced by human walking motion.
To further investigate fine-grained dynamics that are not visible at the longer time scales considered above, we focus on the short-time phase evolution of the residual CIR. Fig. 11 shows a zoomed-in view of the phase $\arg\tilde{h}($20\text{\,}\mathrm{ns}$,t)$, restricted to a duration of $0.1\text{\,}\mathrm{s}$.
As observed in Fig. 11, the phase advanced by approximately $2\pi$ every $T=$33\text{\,}\mathrm{ms}$\$, exhibiting a periodic phase evolution. This indicated that, over this short interval, the residual CIR at $\tau=$20\text{\,}\mathrm{ns}$\$ was dominated by a single complex sinusoidal component with period $T$. Accordingly, $\tilde{h}($20\text{\,}\mathrm{ns}$,t)$ can be locally approximated as $\tilde{h}($20\text{\,}\mathrm{ns}$,t)\propto\mathrm{e}^{\mathrm{j}2\pi\nu t}$, where the frequency $\nu\coloneqq 1/T=$30\text{\,}\mathrm{Hz}$\$ represents the Doppler-induced frequency offset.
V-D3 Physical Interpretation in Terms of Bistatic Range Rate
This periodic phase evolution admits a direct physical interpretation in terms of bistatic range variation. A phase advance of $2\pi$ corresponds to a reduction of the bistatic range by one wavelength $\lambda$ over the same interval $T$. Defining the bistatic range rate as the time derivative of the bistatic range, $\dot{d}_{\mathrm{Tx\text{--}target\text{--}Rx}},$ the observed phase evolution implies $\dot{d}_{\mathrm{Tx\text{--}target\text{--}Rx}}=-\lambda/T.$ This relationship is consistent with the well-known bistatic Doppler relationship [32]
$\nu=-\dot{d}_{\mathrm{Tx\text{--}target\text{--}Rx}}/\lambda.$ (12)
In the present example, the positive phase slope indicates a decreasing bistatic range, corresponding to the target approaching the Tx–Rx pair. Conversely, when the target recedes, the phase slope—and hence the Doppler frequency $\nu$—becomes negative. Therefore, the sign of $\nu$ directly reveals the direction of motion.
Estimating the Doppler frequency $\nu$ thus corresponds to estimating the frequency of the underlying complex sinusoidal component ${\rm e}^{{\rm j}2\pi\nu t}$ discussed above. This estimation is obtained via a delay–Doppler representation of the CIR through Fourier analysis, which generalizes the above short-time analysis, as described in Section VI.
Such fine-grained, phase-coherent delay-domain observations have traditionally required dedicated channel sounders or radar systems. In contrast, the present results demonstrate that this regime can be reached using low-cost COTS Wi-Fi devices. This shift is enabled by the recent availability of wideband Wi-Fi signals with bandwidths of up to $160\text{\,}\mathrm{MHz}$, together with careful preprocessing that allows sufficiently accurate CIR estimation.
V-D4 Comparison with Magnitude-Based Wi-Fi Sensing
Figure 12: Example of a Wi-Fi sensing signal: $\lvert H_{\mathrm{raw}}(100\varDelta f,t)\rvert$.
To clarify the advantage of this phase-coherent interpretation, we now contrast it with conventional magnitude-based Wi-Fi sensing. In such approaches, temporal analysis is performed on the magnitude of the CFR, $\lvert H_{\mathrm{raw}}(f,t)\rvert$. Since the CFR consists of multiple subcarriers, dimensionality reduction is typically applied prior to temporal analysis, for example by selecting representative subcarriers (e.g., [23]) or by applying PCA (e.g., [24]).
As a representative example, Fig. 12 shows $\lvert H_{\mathrm{raw}}(f,t)\rvert$ at a representative subcarrier ($k=100$) over the same interval as Fig. 11. A periodic fluctuation with a period of approximately 33 $3\text{\,}\mathrm{ms}$ was also observed.
Unlike the phase evolution, however, these magnitude fluctuations arise from time-varying interference between human-induced scattering and static multipath components and do not preserve carrier-phase information. Consequently, while the fluctuation rate reflects motion speed, the Doppler sign—and hence whether the target is approaching or receding—cannot be inferred.
V-E Observation of Respiration-Induced Body-Surface Motion
V-E1 Experimental Setup and Range–Time Representation
[…]
Figure 13: Residual CIR $\tilde{h}(\tau,t)$ in a respiration experiment.
Fig. 13LABEL:sub@fig:phase_tau_t presents a delay–time representation of the residual CIR $\tilde{h}(\tau,t)$. In this visualization, the phase of the residual CIR is shown by color, while the magnitude is represented by brightness on a logarithmic scale. A clear periodic pattern was observed around the geometry-based bistatic delay of $20\text{\,}\mathrm{ns}$, suggesting respiration-induced motion of the human target. Note that this response exhibited a noticeable spread in the delay domain. This spread arises from the finite delay resolution, $1/B=$6.3\text{\,}\mathrm{ns}$\$.
V-E2 Phase Evolution and Bistatic Range Interpretation
The phase at $\tau=$20\text{\,}\mathrm{ns}$\$ was further shown in Fig. 13LABEL:sub@fig:phase_t. A clear periodic phase variation with a period of approximately $4\text{\,}\mathrm{s}$ was observed, consistent with the controlled respiration rate of $15$ breaths per minute. The peak-to-peak phase excursion was approximately $3\pi/2$.
To interpret this phase variation to physical motion, let $\varDelta d_{\text{Tx--target--Rx}}$ denote the respiration-induced variation in the bistatic range $d_{\text{Tx--target--Rx}}$. From the relationship $(2\pi/\lambda)\,\varDelta d_{\text{Tx--target--Rx}}=3\pi/2$, the corresponding path-length variation is estimated as $\varDelta d_{\text{Tx--target--Rx}}=3\lambda/4=$40\text{\,}\mathrm{mm}$\$. This value was physically reasonable for respiration-induced body-surface motion. Note that this estimated displacement does not necessarily correspond to a specific anatomical point, but rather reflects respiration-induced variations of dominant reflecting surfaces on the human body.
V-E3 Comparison with Magnitude-Based Wi-Fi Sensing
It is also worth highlighting the difference between the proposed approach and commonly adopted Wi-Fi sensing methods. In many existing studies [23, 33, 34], sensing is performed on the magnitude of CFR, $\lvert H_{\mathrm{raw}}(f,t)\rvert$. Since the CFR spans multiple subcarriers, dimensionality reduction—such as subcarrier selection or PCA—is first applied, and the respiration rate is then estimated by performing an FFT along the time axis of the resulting magnitude time series.
Figure 14: Representative examples of magnitude-based Wi-Fi sensing signals.
As representative examples, the time series of $\lvert H_{\mathrm{raw}}(f,t)\rvert$ at a representative subcarrier ($k=100$) and the PC2 score extracted from $\lvert H_{\mathrm{raw}}(f,t)\rvert$ are shown in Fig. 14. A clear periodic fluctuation was observed, indicating that respiration-induced motion is captured. In this sense, applying an FFT to these signals is sufficient for estimating the respiration rate, and such processing has been widely adopted in existing Wi-Fi sensing studies.
However, compared with Fig. 13LABEL:sub@fig:phase_t, these results exhibited a qualitatively different waveform with substantially larger noise-like fluctuations in magnitude. This is because these results are based on the magnitude of the CSI, which inherently mixes respiration-induced motion with other sources of magnitude variation. Moreover, magnitude-based processing does not preserve the phase corresponding to the bistatic range itself and therefore lacks a direct physical interpretation. As a result, the component associated with respiration is often identified heuristically and varies significantly across devices and environments. In particular, PC1 frequently reflects dominant static or hardware-related variations rather than motion-induced dynamics, as reported in prior studies [24, 23].
In contrast, the LoSRef-based method exploits the temporal evolution of the phase, enabling a direct physical interpretation in terms of bistatic range variations. Unlike magnitude-based sensing, this phase-based approach preserves propagation-induced information and is less sensitive to AGC-induced gain fluctuations.
VI Delay–Doppler Response Estimation of Moving Human Target
This section investigates delay–Doppler response estimation of a moving human target using phase-coherent Wi-Fi CIRs. To this end, we first characterize the impact of irregular temporal sampling in practical Wi-Fi systems and introduce time-domain resampling. We then construct a delay–Doppler representation and demonstrate range–Doppler estimation through experiments.
VI-A Inter-Packet Timing Variations in Contention-Based Wi-Fi
Although frame transmissions are nominally scheduled at regular intervals, the actual observation times at the Rx are not guaranteed to be uniformly spaced. This is due to contention-based medium access, occasional missed receptions, and the removal of outliers.
Figure 15: Temporal variation and empirical distribution of the observed inter-packet intervals. Although the nominal sampling interval is $1\text{\,}\mathrm{ms}$, the intervals exhibit noticeable fluctuations, with a median of $1.069\text{\,}\mathrm{ms}$.
This timing variability was experimentally examined in Fig. 15. Fig. 15LABEL:sub@fig:intersample_time illustrates the inter-packet intervals computed from the observed measurements in the experiment described in Section V-D, where packet transmissions were intended to follow a $1\text{\,}\mathrm{ms}$ interval. The inter-packet intervals were spread over a range from approximately $0.5\text{\,}\mathrm{ms}$ to $6.5\text{\,}\mathrm{ms}$, indicating significant deviations from the nominal sampling period.
Fig. 15LABEL:sub@fig:intersample_time_cdf shows the corresponding empirical cumulative distribution function (CDF) of the inter-packet intervals. The observed intervals exhibited a median of $1.069\text{\,}\mathrm{ms}$, with a median absolute deviation of $0.017\text{\,}\mathrm{ms}$, indicating small but non-negligible timing jitter around the nominal sampling interval.
VI-B Uniform Temporal Resampling for Phase-Coherent Doppler Estimation
Since Doppler estimation generally requires uniformly spaced observations in time, a uniform time grid is constructed, where the grid spacing $\varDelta t$ is set to the median inter-packet interval, i.e., $\varDelta t=$1.069\text{\,}\mathrm{ms}$\$ in the present measurements. Each observed CIR snapshot is first associated with its temporally nearest grid point. The remaining unassociated grid points are then filled by interpolation: the magnitude is interpolated linearly, while the phase is interpolated after phase unwrapping so as to preserve temporal continuity. Although more advanced treatments of non-uniform sampling are possible, the above procedure is sufficient for accurate phase-coherent Doppler estimation in this work. Although more advanced treatments of non-uniform sampling are possible, the above grid-based resampling procedure is sufficient for accurate Doppler estimation in this work.
VI-C Delay–Doppler Representation and Doppler Interpretation
The delay–Doppler representation is obtained by Fourier transforming the CIR $h(\tau,t)$ with respect to time $t$, as is well established in the literature [28, 27]. While Section V-D illustrated Doppler behavior over a short time interval at a fixed delay, the delay–Doppler representation generalizes this observation by characterizing Doppler frequency as a function of delay. In practice, we apply an STFT along the time axis to the phase-coherent and uniformly resampled CIR $h(\tau,t)$, yielding a time-varying delay–Doppler representation $s(\tau,\nu;t)$, where $t$ denotes the center time of the analysis window. [^2] The CIR $h(\tau,t)$ may be taken either before or after clutter removal in Section IV-C, since clutter components are suppressed on a per-STFT-segment basis. This representation jointly captures the temporal evolution of multipath delays and their associated Doppler frequencies.
Although the experiment employs a bistatic configuration, the Tx–Rx separation is sufficiently small that the geometry can be approximated as monostatic. In this limit, the bistatic range $d_{\text{Tx--target--Rx}}$ reduces to twice the one-way Tx/Rx–target distance, and the bistatic range rate becomes
$\dot{d}_{\text{Tx--target--Rx}}=2v,$ (13)
where $v$ denotes the radial velocity defined as the time derivative of the one-way Tx/Rx–target distance. Substituting (13) into (12) yields
$\nu=-{2v}/{\lambda}.$ (14)
For ease of interpretation, we therefore employ an effective radial velocity scaled as $\nu\lambda/2$ in the following discussion, instead of the bistatic range rate.
VI-D Experimental Results: Range–Doppler Estimation
TABLE III: STFT Parameters for Doppler Analysis
STFT segment length $256$
STFT overlap $224$
Doppler-domain interpolation rate 88
STFT window Hann
The main STFT parameters are summarized in Table III. Similar to the Fourier transform from the frequency domain to the delay domain in (3), a Hann window was applied to each STFT segment of length $256$ samples to suppress spectral leakage, with an overlap of $224$ samples (87.587.5%), corresponding to a hop size of $32\text{\,}\mathrm{ms}$. With an STFT segment of $256$ samples and a time grid of $\varDelta t\approx$1\text{\,}\mathrm{ms}$\$, the resulting Doppler frequency resolution, determined by the effective STFT window duration, is approximately $\varDelta\nu\approx$3.9\text{\,}\mathrm{Hz}$\$, which is sufficient to resolve human motion-induced Doppler components. In addition, Doppler-domain oversampling with a factor of $8$ was performed by zero-padding in the time domain. Furthermore, the mean value of each STFT segment was removed to mitigate DC components. This operation can be interpreted as applying clutter removal described in Section IV-C at the level of individual STFT segments.
Figure 16: Representative snapshots of ${s}(\tau,\nu;t)$. As the target passes near the Tx–Rx pair, the Doppler frequency changes sign from positive to negative, and the propagation delay reaches a minimum around this passage.
A representative example of the range–Doppler estimation ${s}(\tau,\nu;t)$, obtained from the moving human target measurements described in Section V-D, was shown in Fig. 16. The evolution of the delay–Doppler structure over time clearly reflected the target motion relative to the Tx–Rx pair.
Figure 17: Estimated bistatic range $c\tau^{\star}$ and Doppler frequency $\nu^{\star}$ of a moving human target, extracted by selecting the maximum of $s(\tau,\nu;t)$ at each time $t$. The curve indicates the geometry-based value derived from the motion-tracking system, shown for reference.
Fig. 17 shows the estimated bistatic range $c\tau^{\star}$ and Doppler frequency $\nu^{\star}$ of the moving human target, together with the corresponding geometry-based values derived from the motion-tracking system. Here, $\tau^{\star}$ and $\nu^{\star}$ were obtained by selecting, at each time instant $t$, the maximum of the delay–Doppler representation $s(\tau,\nu;t)$, i.e.,
$(\tau^{\star},\nu^{\star})\coloneqq\arg\max_{(\tau,\nu)}s(\tau,\nu;t).$ (15)
Overall, the estimated bistatic range and Doppler frequency exhibited trends that are broadly consistent with the geometry-based reference. It should be noted that the tracking system captures the motion of a marker attached to the target’s head, while the Doppler estimates may be influenced by distributed scattering from multiple body parts. Therefore, perfect agreement is not expected.
More specifically, Fig. 17LABEL:sub@fig:range-time-peak corresponds to the delay–time heatmap shown in Fig. 10LABEL:sub@fig:h_tau-t. The bistatic range shown here is given by $c\tau^{\star}$ extracted from the delay–Doppler representation, rather than by directly taking the peak of the residual CIR $\tilde{h}(\tau,t)$.
In Fig. 17LABEL:sub@fig:doppler-time-peak, the estimated Doppler frequency $\nu^{\star}$ is shown together with the corresponding effective radial velocity $\nu^{\star}\lambda/2$, evaluated using (14), on the right axis. For typical Wi-Fi carrier frequencies, human motion gives rise to Doppler frequencies on the order of several tens of hertz [24, 35]. With a uniform time grid of $\varDelta t\approx$1\text{\,}\mathrm{ms}$\$, the resulting effective Doppler bandwidth of $\pm$500\text{\,}\mathrm{Hz}$\$ is therefore sufficient for capturing such motion-induced Doppler components.
Figure 18: Doppler–time representation obtained by incoherent integration of the delay–Doppler response ${s}(\tau,\nu;t)$.
The peak of $s(\tau,\nu;t)$, i.e., $(\tau^{\star},\nu^{\star})$, provides intuitive point estimates of the bistatic range and Doppler frequency, but does not capture the full Doppler energy distribution. To complement this point-wise representation, we additionally examined a Doppler–time representation obtained by incoherently integrating the power $\lvert s(\tau,\nu;t)\rvert^{2}$ along the delay axis, as shown in Fig. 18. This representation visualizes the temporal Doppler energy and exhibits a similar trend to that in Fig. 17LABEL:sub@fig:doppler-time-peak; however, the two representations are not directly comparable, as they are derived through different aggregation procedures.
VI-E Comparison with Frequency-Domain Magnitude-Based Wi-Fi Sensing
For comparison, we also present a representative conventional Wi-Fi sensing result that operates purely in the frequency domain and does not rely on the CIR representation. As discussed in Section II-A, many existing approaches perform sensing directly on the magnitude of the CFR, $\lvert H_{\mathrm{raw}}(f,t)\rvert$. In this paper, PCA is used as a representative example of dimensionality reduction, as in [24].
Figure 19: Doppler-related time–frequency representation obtained from the STFT of the PC3 score of $\lvert H_{\mathrm{raw}}(f,t)\rvert$.
Among the resulting PC scores, the STFT is applied to the PC3 score, which exhibits clear target-induced temporal variations. It should be noted, however, that the PC capturing motion-induced dynamics is device- and environment-dependent [24, 23], and its identification often relies on empirical or heuristic criteria rather than a physical model. The resulting Doppler-related time–frequency representation is shown in Fig. 19.
A key limitation of this magnitude-based frequency-domain approach is that it does not yield a physically meaningful Doppler frequency, because the underlying propagation-induced phase evolution is not preserved, as explained in Section II-A with reference to [26]. As a result, the spectrum in Fig. 19 is symmetric with respect to frequency sign and effectively collapses positive and negative Doppler components, making it impossible to distinguish approaching and receding motions.
By contrast, the proposed LoSRef-based approach operates on the complex-valued residual CIR, whose Doppler spectrum is inherently asymmetric with respect to frequency sign. This asymmetry preserves Doppler sign information and enables direct discrimination between approaching and receding motions, which is fundamentally impossible with magnitude-based processing.
VII Discussion
The experimental validation was performed using the Intel AX210 chipset. While multiple tools for CSI extraction from IEEE 802.11ax devices exist [36, 37], they are all restricted to the Intel AX2xx family, precluding validation on alternative chipsets. Nevertheless, the LoSRef-based framework is chipset-agnostic, as the subsequent processing after phase-coherent CIR construction follows standard bistatic radar signal processing principles, and is therefore applicable to any system that provides a complex-valued CFR and a CIR with sufficient delay resolution.
Recent studies have shown that range–Doppler information can be extracted from CSI using COTS IEEE 802.11ax Wi-Fi devices under monostatic configurations [38]. However, extending such capabilities to bistatic scenarios remains challenging, primarily due to the lack of a reliable reference signal between the Tx and Rx.
VIII Conclusion
This paper presented a LoSRef-based bistatic Wi-Fi radar framework that enables phase-coherent delay–Doppler analysis using only unmodified COTS Wi-Fi devices. LoSRef leverages the Tx–Rx LoS path as a delay and phase reference; its inherent temporal stability enables consistent phase-coherent alignment across successive packet receptions. As a result, the proposed framework eliminates the need for wired references or dedicated reference antennas and achieves phase-coherent bistatic radar operation with a single-antenna, single-port Rx, thereby breaking a key deployment barrier in existing Wi-Fi radar systems.
Through representative human gait and respiration experiments, we confirmed that phase-coherent bistatic radar functionality can be realized within typically deployed, communication-oriented Wi-Fi systems. Specifically, the proposed framework enables physically meaningful delay and Doppler estimates, including gait direction discrimination and respiration-induced sub-wavelength displacement measurements.
The proposed LoSRef-based framework further suggests that Wi-Fi sensing can move beyond application-specific processing. Once phase-coherent representations are restored, radar signal processing techniques can be directly applied to Wi-Fi sensing, pointing to a fruitful research direction for future ISAC system design.
Acknowledgment
The author would like to thank Mr. Genichiro Shiino for providing the experimental data used in this study and for valuable discussions throughout the course of this work. The author also thanks Prof. Katsuyuki Haneda for his insightful comments on an early draft of this paper, and Prof. Hiromitsu Awano for his valuable comments on vital sensing.